OpenStudy (anonymous):

Graph the line (5,6) with slope -1/2

OpenStudy (haseeb96):

(y-y1)=m(x-x1) since m=-1/2 (x1,y1)=(5,6) now just put it in the above formula and get the equation of a line

OpenStudy (haseeb96):

y-6=-1/2 (x-5) -2y+12=x-5 x+2y-5-12=0 x+2y -17=0 this is the equation of line

OpenStudy (anonymous):

Ho do i graph it"?

OpenStudy (anonymous):

Haseeb96

OpenStudy (haseeb96):

do u know how to draw a graph of an equation ax+by+c=0 ?